OpenStudy (anonymous):

Write the word sentence as an inequality A number no less than -4 A number b multiplied by -5 is at most -3/4

OpenStudy (anonymous):

\[b \ge -4\] \[b(-5) \le -\frac{ 3 }{ 4 }\]